# Semantic Aware Linear Transfer by Recycling Pre-trained Language Models for Cross-lingual Transfer

### v1

@@ Abstract
Seungyoon Lee, Seongtae Hong, Hyeonseok Moon, Heuiseok Lim Korea University, Republic of Korea {dltmddbs100, ghdchlwls123, glee889, limhseok}@korea.ac.kr
@@ 1 Introduction
To address this challenge, a promising line of studies has focused on cross-lingual transfer Artetxe et al. (2020); Tran (2020); Gee et al. (2022); Ostendorff and Rehm (2023); Dobler and De Melo (2023); Remy et al. (2024); Liu et al. (2024); Mundra et al. (2024); Ye et al. (2024). These efforts include replacing a source vocabulary with a target language vocabulary, simply initializing embeddings with newly manipulated embeddings stem from source model Minixhofer et al. (2022); Dobler and De Melo (2023); Liu et al. (2024); Mundra et al. (2024). However, newly initialized embedding that relies solely on the source model inevitably experiences limited expressiveness in the target language due to the immature learning process of the source model that does not mainly concentrate on the target language. Moreover, most approaches adopt small encoder architecture such as BERT-like models Devlin et al. (2019); Liu (2019); Conneau et al. (2020) as a source model, leaving their applicability on contemporary decoder-based LLMs unexplored Gogoulou et al. (2022); Gee et al. (2022); Zeng et al. (2023); Dobler and De Melo (2023); Ye et al. (2024).
In this work, we propose Semantic Aware Linear Transfer (SALT), a novel cross-lingual transfer method that leverages the rich representational power of target language embeddings in traditional Pre-trained Language Models (PLMs) to convert English-centric LLMs into target language–specialized LLMs. To facilitate the transfer, we identify the semantically closest tokens for each non-shared token from a shared vocabulary space. Based on these semantically similar tokens, we fit a unique linear regression to transfer the embeddings from the PLM’s space to the LLM’s space. This procedure enables non-shared vocabulary to be transferred to LLM space while retaining semantic representation embedded in PLM embedding.
@@ 2 Related Work
Figure 1: Summary of SALT. By using the paired embeddings of semantically similar tokens for each non-shared token, we create a unique least square matrix to transfer from PLM to LLM.
@@
However, initializing embedding relying on the source model’s embedding inherently lacks crucial information specialized for the target language since the source embedding does not prioritize target language adaptation. Given this concern, we adopt embeddings from PLMs dedicated to the target language to convey richer semantic features that the source model fails to capture during the transfer Rajaee and Pilehvar (2022). Based on this approach, we propose a new cross-lingual transfer method that recycles target language PLMs to elicit LLMs for better adaptation in target languages.
@@ 3 SALT
For each $V_{s}$ and $V_{t}$, we use external static embeddings from fastText Bojanowski et al. (2017), which are trained on the target languages, to extract auxiliary embeddings. We select fastText trained in multiple languages, as it can provide embeddings of various words using combinations of multiple subwords. For rare tokens that do not exist in fastText or cannot be formed via subword combinations, we initialize them from normal distribution with mean and standard deviation from $\boldsymbol{E}_{s}$.
@@
Our primary goal is to transform non-shared vocabulary embeddings into the source model space by leveraging semantically similar shared tokens’ paired embeddings. We employ the cheap and fast Least Squares Transform approach. To this end, we stack the source and target embeddings of the selected nearest tokens, which is $\boldsymbol{E}^{\prime}_{t_{i}}$ and $\boldsymbol{E}^{\prime}_{s_{i}}$ respectively. Note that selected tokens have embeddings from $\boldsymbol{E}_{s}$ and $\boldsymbol{E}_{t}$ since we only extract the nearest tokens from $V_{shared}$.
@@
As a final step, we perform additional training with unlabeled target language corpus, called Language Adaptive Continual Pre-training, as a final step to align the weights between transferred embedding and source model layers in the target language. This stage is crucial after the initialization or transformation of embedding since new embedding lacks alignment with the upper layers. Thus, this stage is considered essential as a post-transfer task Chau et al. (2020); Dobler and De Melo (2023); Liu et al. (2024); Mundra et al. (2024).
@@ 4.1 Models
For the target model, we select PLMs with published references of various architectures for each language to ensure reliability, focusing exclusively on base-level models. By considering these widely used standard PLMs, we ensure the generalizability of our proposed method. To be specific, we use PLMs with BERT Devlin et al. (2019), GPT Brown et al. (2020), and T5 Raffel et al. (2020) architectures for each language to verify the scalability. Selected target models are provided in Appendix B.
@@ 4.2 Baseline
The One For All (OFA) Framework Liu et al. (2024) uses multilingual static word vectors to inject alignment knowledge into the new subword embeddings. With the embedding factorization approach via Singular Value Decomposition, OFA initializes embedding by leveraging a weighted average of the source embedding.
@@ 4.4 Evaluation
Table 1: Performance on knowledge-based benchmarks. HS and TQA stand for HellaSwag and TruthfulQA. We bold the best result in each section under the same condition. Results correspond to decoder-based PLMs.
@@ 5.3 Effects In Language Understanding
In general, most PLMs mainly focusing on the target language often consider both English and the target language in the training stage, creating a shared space for both. By projecting this fused space to new embedding, SALT successfully blends the source model’s English ability with newly initialized embeddings for the target language, leading to notable benefits in cross-lingual understanding.
@@ 5.4 Extension to Target Models with Different Architectures
We also compare performance using target models whose architectures differ from the source model (e.g., BERT and T5). Their embedding formations differ based on learning objectives. We find that SALT can be a valid transfer method for target models with different structures from the source model.
As illustrated in Table 4, the decoder-based target model mostly achieves the best average results. However, employing PLMs with different structures as the target model can perform better in some cases. In general, T5 shows robust performance. For German, using T5 generally outperforms GPT in both source models. In addition, for German and Arabic, BERT’s performance gap with GPT remains within about one point on average. This confirms that the target model does not necessarily need to share the same structure as the source model. Although target models are trained on different learning objectives, transferring their embeddings to the source LLMs still provides sufficient expressiveness for the target language. Given that various kinds of PLMs exist, this demonstrates the flexibility in that our method is not restricted to decoder-type target models and can be extended to various architectures.
@@ 6 Conclusion
We propose SALT, a novel cross-lingual transfer method that recycles the embeddings of PLMs for transferring from English-centric LLMs to target language-specific LLMs. By deriving a unique linear regression line that takes into account the semantic similarity for each non-overlapping vocabulary embedding, we map the rich semantic features in PLM’s embedding onto the LLM’s embedding. A series of experiments reveal that SALT benefits from the rich expressiveness of PLM’s embedding in the target language. Also, SALT can promote better initialization and faster convergence for language adaptation, enhancing generalization to the target language. Notably, among existing initialization methods, SALT preserves English capabilities more effectively and demonstrates greater superiority in cross-lingual setups. We also investigate expandability by experimenting with multiple types of target models, indicating that SALT can be applied to various PLM architectures. For future work, we plan to investigate methods that facilitate the transfer of target language-specific knowledge embedded within model layers.
@@
In our study, our methodology is based on PLMs, making the presence of PLMs essential. Our selection of languages in this paper is based on this consideration. For our experiments in diverse types of PLMs, low-resource languages where these PLMs are unavailable are out of our scope. Given that SALT benefits from target language PLMs’ rich representation ability, SALT is likely to be an even more effective method in low-resource languages if PLMs exist.
In the evaluation, we did not investigate the impact of instruction tuning. Our focus lies in leveraging foundational models for transfer and evaluating the benefits derived from PLM embeddings and the linguistic ability of the transferred model since our approach is only related to the initialization stage. Our transfer method provides a robust starting point for embeddings, laying the groundwork for future instruction tuning to build models that perform better in the target language.
Also, we evaluate SALT only for models that have large vocabulary sizes accounting for a relatively heavy portion of embedding in the model’s total parameters. Given that one of the objectives of cross-lingual transfer is mitigating overload caused by irrelevant tokens to the target language and pursuing effectiveness, this aligns with the goal. In principle, SALT can also be applied to larger models (>7b), and language-specific LLMs can be used for target models instead of PLMs. However, our focus is on the usability of the previous PLMs and we would call for exploration in this direction.
@@ Appendix B Selected Target Models
Appendix C Hyper-parameters and Training Details
@@ Appendix E Length & Parameter Footprint Analysis
Similarly, SALT results in a substantial decrease of parameters—24% in XGLM and 17% in Gemma. The reduction is due to a decrease in embedding parameters by eliminating vocabulary useless in the target language within LLMs. SALT reduces unnecessary loads for the target language and enables the use of a language-specific tokenization strategy that accounts for the unique characteristics of the target language derived from the considerable amount of monolingual corpus.
Model Languages

### v2

@@ Abstract
Seungyoon Lee, Seongtae Hong, Hyeonseok Moon, Heuiseok Limthanks: Corresponding authors Korea University, Republic of Korea {dltmddbs100, ghdchlwls123, glee889, limhseok}@korea.ac.kr
@@ 1 Introduction
To address this challenge, a promising line of studies has focused on cross-lingual transfer Artetxe et al. (2020); Tran (2020); Gee et al. (2022); Dobler and De Melo (2023); Remy et al. (2024); Liu et al. (2024); Mundra et al. (2024); Ye et al. (2024). These efforts include replacing a source vocabulary with a target language vocabulary, simply initializing embeddings with newly manipulated embeddings stem from source model Minixhofer et al. (2022); Dobler and De Melo (2023); Liu et al. (2024); Mundra et al. (2024).
However, a newly initialized embedding that relies solely on the source model inevitably experiences limited expressiveness in the target language due to the immature learning process of the source model, which does not mainly concentrate on the target language. Moreover, most approaches adopt small encoder architectures such as BERT-like models Devlin et al. (2019); Liu (2019); Conneau et al. (2020) as a source model, leaving their applicability on contemporary decoder-based LLMs unexplored Gogoulou et al. (2022); Gee et al. (2022); Zeng et al. (2023); Dobler and De Melo (2023); Ye et al. (2024).
In this work, we propose Semantic Aware Linear Transfer (SALT), a novel cross-lingual transfer method that leverages the rich representational power of target language embeddings in traditional Pre-trained Language Models (PLMs) [^1] In this work, we define PLMs to small-scale language models with only millions of parameters, developed prior to the advent of LLMs. to convert English-centric LLMs into target language–specialized LLMs. To facilitate the transfer, we identify the semantically closest tokens for each non-shared token from a shared vocabulary space. Based on these semantically similar tokens, we fit a unique linear regression to transfer the embeddings from the PLM’s space to the LLM’s space. This procedure enables non-shared vocabulary to be transferred to LLM space while retaining the semantic representation embedded in PLM embedding.
@@ 2 Related Work
Figure 1: Summary of SALT. By using the paired embeddings of semantically similar tokens for each non-shared token, we create a unique least square matrix $X_{t_{i}}$ to transfer from PLM to LLM. $k$ denotes the number of selected nearest tokens among overlapping tokens for linear transfer, and $h_{t}$ and $h_{s}$ refer to the hidden dimensions of the target and source models.
@@
However, initializing embedding relying on the source model’s embedding inherently lacks crucial information specialized for the target language since the source embedding does not prioritize target language adaptation. Given this concern, we adopt embeddings from PLMs dedicated to the target language to convey richer semantic features that the source model fails to capture during the transfer. Based on this approach, we propose a new cross-lingual transfer method that recycles target language PLMs to elicit LLMs for better adaptation in target languages.
@@ 3 SALT
For each $V_{s}$ and $V_{t}$, we use external static embeddings from fastText Bojanowski et al. (2017), which are trained on the target languages, to extract auxiliary embeddings. We select fastText trained in multiple languages, as it can provide embeddings of various words using combinations of multiple subwords. For rare tokens that do not exist in fastText or cannot be formed via subword combinations, we initialize them from a normal distribution with mean and standard deviation from $\boldsymbol{E}_{s}$.
@@
Our primary goal is to transform non-shared vocabulary embeddings into the source model space by leveraging semantically similar shared tokens’ paired embeddings. We employ the cheap and fast Least Squares Transform approach. To this end, we stack the source and target embeddings of the selected nearest tokens, which are $\boldsymbol{E}^{\prime}_{t_{i}}$ and $\boldsymbol{E}^{\prime}_{s_{i}}$ respectively. Note that selected tokens have embeddings from $\boldsymbol{E}_{s}$ and $\boldsymbol{E}_{t}$ since we only extract the nearest tokens from $V_{shared}$.
@@
As a final step, we perform additional training with an unlabeled target language corpus, called Language Adaptive Continual Pre-training, as a final step to align the weights between transferred embedding and source model layers in the target language. This stage is crucial after the initialization or transformation of the embedding since the new embedding lacks alignment with the upper layers. Thus, this stage is considered essential as a post-transfer task Chau et al. (2020); Dobler and De Melo (2023); Liu et al. (2024); Mundra et al. (2024).
@@ 4.1 Models
For the target model, we select PLMs with published references of various architectures for each language to ensure reliability, focusing exclusively on base-level models. By considering these widely used standard PLMs, we ensure the generalizability of our proposed method. To be specific, we use PLMs with BERT Devlin et al. (2019), GPT Brown et al. (2020), and T5 Raffel et al. (2020) architectures for each language to verify the scalability. Unless otherwise specified, our experiments use decoder-based PLMs (GPT) as the source model. Selected target models are provided in Appendix B.
@@ 4.2 Baseline
The One For All (OFA) Framework Liu et al. (2024) uses multilingual static word vectors to inject alignment knowledge into the new subword embeddings. With the embedding factorization approach via Singular Value Decomposition, OFA initializes the embedding by leveraging a weighted average of the source embedding.
@@ 4.4 Evaluation
Table 1: Performance on knowledge-based benchmarks. HS and TQA stand for HellaSwag and TruthfulQA. We bold the best result in each section under the same condition.
@@ 5.3 Effects In Language Understanding
In general, most PLMs mainly focusing on the target language often consider both English and the target language in the training stage, creating a shared space for both. By projecting this fused space to a new embedding, SALT successfully blends the source model’s English ability with newly initialized embeddings for the target language, leading to notable benefits in cross-lingual understanding.
@@ 5.4 Extension to Target Models with Different Architectures
To demonstrate the scalability of SALT, we compare performance using target models whose architectures differ from the source model (e.g., BERT and T5). Their embedding formations differ based on learning objectives. We find that SALT can be a valid transfer method for target models with different structures from the source model.
As illustrated in Table 4, the decoder-based target model mostly achieves the best average results. However, employing PLMs with different structures as the target model performs better in some cases. For German, using T5 generally outperforms GPT in both source models. In addition, for German and Arabic, BERT’s performance gap with GPT remains within about one point on average.
We also report the evaluation loss according to various types of target models in SALT in Table 5. In some cases, BERT achieves lower evaluation loss compared to GPT. Although learning patterns may vary depending on the underlying target model, these observations suggest that other types of models can be integrated into decoder-only LLMs from a transfer standpoint. Given that various kinds of PLMs exist, SALT can be a flexible method that is not restricted to decoder-type target models and can be extended to various architectures.
Source Target German Arabic Vietnamese
30% 60% 90% 30% 60% 90% 30% 60% 90%
XGLM BERT 3.89 3.82 3.80 3.76 3.25 3.24 3.49 3.40 3.40
GPT 4.08 4.00 3.99 3.83 3.68 3.67 3.25 3.18 3.17
T5 3.77 3.71 3.70 4.27 4.14 4.13 3.07 3.02 3.01
Gemma BERT 2.81 2.77 2.73 2.39 2.24 2.22 2.50 2.45 2.37
GPT 3.09 2.88 2.81 2.76 2.54 2.51 2.37 2.32 2.24
T5 3.07 2.90 2.85 3.15 2.91 2.86 2.32 2.29 2.22
Table 5: Evaluation loss at various percentages of total steps for each target model in SALT.
@@ 6 Conclusion
We propose SALT, a novel cross-lingual transfer method that recycles the embeddings of PLMs for transferring from English-centric LLMs to target language-specific LLMs. By deriving a unique linear regression line that takes into account the semantic similarity for each non-overlapping vocabulary embedding, we map the rich semantic features in PLM’s embedding onto the LLM’s embedding. A series of experiments reveals that SALT benefits from the rich expressiveness of PLM’s embedding in the target language. Also, SALT can promote better initialization and faster convergence for language adaptation, enhancing generalization to the target language. Notably, among existing initialization methods, SALT preserves English capabilities more effectively and demonstrates greater superiority in cross-lingual setups. We also investigate expandability by experimenting with multiple types of target models, indicating that SALT can be applied to various PLM architectures. For future work, we plan to explore methods that facilitate the transfer of target language-specific knowledge embedded within model layers.
@@
Our methodology is based on target language-specialized PLMs, making their presence essential. Our selection of languages in this paper is based on this consideration. For our experiments in diverse types of PLMs, low-resource languages where these PLMs are unavailable are out of our scope. Given that SALT benefits from target language PLMs’ rich representation ability, SALT is likely to be an even more effective method in low-resource languages if PLMs exist.
In the evaluation, we did not investigate the impact of instruction tuning. Our focus lies in leveraging foundational models for transfer and evaluating the benefits derived from PLM embeddings and the linguistic ability of the transferred model, since our approach is only related to the initialization stage. Our transfer method provides a robust starting point for embeddings, laying the groundwork for future instruction tuning to build models that perform better in the target language.
Also, we evaluate SALT only for models with large vocabulary sizes accounting for a relatively heavy portion of embedding in the model’s total parameters. Given that one of the objectives of cross-lingual transfer is mitigating overload caused by irrelevant tokens to the target language and pursuing effectiveness, this aligns with the goal. In principle, SALT can also be applied to larger models (>7b), and language-specific LLMs can be used for target models instead of PLMs. However, our focus is on the usability of the previous PLMs and we would call for exploration in this direction.
@@
Acknowledgments
This research was supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2021R1A6A1A03045425). This work was supported by Institute for Information & communications Technology Promotion (IITP) grant funded by the Korea government (MSIT) (RS-2024-00398115, Research on the reliability and coherence of outcomes produced by Generative AI). This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) under the Leading Generative AI Human Resources Development (IITP-2025-R2408111) grant funded by the Korea government (MSIT). This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) under the artificial intelligence star fellowship support program to nurture the best talents (IITP-2025-RS-2025-02304828) grant funded by the Korea government (MSIT).
@@ Appendix A Further Details of SALT
Insufficient overlap potentially hinders the formation of a linear regression line conveying the capabilities of target-language specialized PLMs. Consequently, comparatively lower performance in Arabic than in other languages in our experiments may stem from the low coverage rate of overlapping tokens.
Implementation
We apply the same mechanism to both the embedding layer and the LM head layer. Accordingly, both the embedding layer and the head layer are initialized with the vocabulary of the target language model. If the LM embedding layer has dimensions of (vocab_size × hidden_size), then the head layer adopts dimensions of (hidden_size × vocab_size), which is the transpose of the LM embedding matrix. However, this may vary depending on whether the model uses tied embedding weights. In our case, we used a model where the LM head and the embedding layer share the same weights; therefore, we initialized the new embedding matrix and then assigned its transpose to the LM head. If the weights are not shared, it would be necessary to apply SALT to the LM head weight matrix separately.
@@ Appendix B Selected Target Models
Appendix C Training Details
@@ Appendix E Length & Parameter Footprint Analysis
Similarly, SALT results in a substantial decrease of parameters: 24% in XGLM and 17% in Gemma. The reduction is due to a decrease in embedding parameters by eliminating vocabulary that is unnecessary in the target language within LLMs. SALT reduces unnecessary loads for the target language and enables the use of a language-specific tokenization strategy that accounts for the unique characteristics of the target language derived from a considerable amount of monolingual corpus.
Model Language